OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of -2, 3 of multiplicity 2.Explain how you would verify the zeros of f(x)

OpenStudy (anonymous):

The polynomial has the multiplicity 2 right? Or is it the roots that have the multiplicity?

OpenStudy (anonymous):

The polynomial has the multiplicity 2

OpenStudy (anonymous):

Ok then you know that since the multiplicity is 2, the maximum amount of real roots is 2.

OpenStudy (anonymous):

So yo can just do f(x) = (x+2)(x-3)

OpenStudy (anonymous):

To find the roots of an equation, you set f(x) = 0 and solve. So when we check it, it is 0 = (x+2)(x-3). x=-2 and x=3 will be your roots

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

Yes It Do.

OpenStudy (anonymous):

ok cool. Tell me if you need more clarification

OpenStudy (anonymous):

Okay.Thank You So MUCH

OpenStudy (anonymous):

Can u help me with another one please